# M.A. Basiri & S. Khanmohammadi

### v2

@@ Abstract
The combination of machine learning (ML) and sparsity-promoting techniques is enabling direct extraction of governing equations from data, revolutionizing computational modeling in diverse fields of science and engineering. The discovered dynamical models could be used to address challenges in climate science, neuroscience, ecology, finance, epidemiology, and beyond. However, most existing sparse identification methods for discovering dynamical systems treat the whole system as one without considering the interactions between subsystems. As a result, such models are not able to capture small changes in the emergent system behavior. To address this issue, we developed a new method called Sparse Identification of Nonlinear Dynamical Systems from Graph-structured data (SINDyG), which incorporates the network structure into sparse regression to identify model parameters that explain the underlying network dynamics. We showcase the application of our proposed method using several case studies of neuronal dynamics, where we model the macroscopic oscillation of a population of neurons using the extended Stuart-Landau (SL) equation and utilize the SINDyG method to identify the underlying nonlinear dynamics. Our extensive computational experiments validate the improved accuracy and simplicity of discovered network dynamics when compared to the original SINDy approach. Data-Driven Modeling, Sparse Regression, Network Dynamics, Coupled Oscillators, Neuronal Population
@@ 1 Introduction
Despite the encouraging results by models like SINDy, they are not designed for complex networked systems with many interacting components such as biological neural networks in the brain [15], power grids [18], social networks [1], and traffic systems [24]. In these systems, the state of each node is influenced not only by its own dynamics but also by the interactions with its connected neighbors [22]. This is because complex system behavior arises from the dynamic interactions between the nodes, not just from the nodes themselves [9]. Several methods have been developed to infer network dynamics while taking into account the interaction between nodes [9, 4]. However, majority of these methods typically do not leverage existing knowledge of graph structures or directly incorporate them into the modeling of node-level dynamics.
In this paper, we propose an extension of the SINDy method, termed SINDyG (Sparse Identification of Nonlinear Dynamics for Graph-structured data), which explicitly accounts for the network structure during sparse regression. By incorporating the connections between nodes, our approach uncovers not only the individual dynamics of each node but also the mechanisms by which they interact, leading to more accurate and interpretable models. We test our method using several case studies of neuronal dynamics, where the goal is to identify nonlinear interactions between populations of neurons. Our results indicate improvements to model accuracy and simplicity when compared to the original SINDy method.
The rest of the work is organized as follows: In Sec. II, we discuss the Sparse Identification of Nonlinear Dynamics (SINDy) [3] and all the necessary information about graph-structured data. In Sec. III, we explain our proposed method for data-driven modeling of complex and interconnected dynamical systems. Next, Section IV presents the results of our method using several case studies based on neuronal dynamics. Finally, we conclude the paper in Section V.
@@ 2 Preliminaries
The following sections provide a description of Sparse Identification of Nonlinear Dynamics (SINDy) and Graph-structured time series data, which are going to be frequently used in this paper.
@@ 2.1 Sparse Identification of Nonlinear Dynamics (SINDy)
$\frac{\mathrm{d}\mathbf{x}(t)}{\mathrm{d}t}=\mathbf{F}\left(\mathbf{x}(t)\right)$ (1)
@@
In order to compute the function $\mathbf{F}$ from data, a time series of system’s state variables $\mathbf{x}(t)$ is collected. Also, the derivatives of state variables $\mathbf{\dot{x}}(t)$ are measured or approximated numerically from $\mathbf{x}(t)$ and arranged in the following matrix:
$\dot{\mathbf{X}}=\left[\begin{array}[]{c}\dot{\mathbf{x}}^{T}\left(t_{1}\right)\\ \dot{\mathbf{x}}^{T}\left(t_{2}\right)\\ \vdots\\ \dot{\mathbf{x}}^{T}\left(t_{T}\right)\end{array}\right]=\left[\begin{array}[]{cccc}\dot{x}_{1}\left(t_{1}\right)&\dot{x}_{2}\left(t_{1}\right)&\cdots&\dot{x}_{K}\left(t_{1}\right)\\ \dot{x}_{1}\left(t_{2}\right)&\dot{x}_{2}\left(t_{2}\right)&\cdots&\dot{x}_{K}\left(t_{2}\right)\\ \vdots&\vdots&\ddots&\vdots\\ \dot{x}_{1}\left(t_{T}\right)&\dot{x}_{2}\left(t_{T}\right)&\cdots&\dot{x}_{K}\left(t_{T}\right)\end{array}\right]$
@@
$\mathbf{\Theta}(\mathbf{X})=\left[\begin{array}[]{cccccccc}\mid&\mid&\mid&\mid&&\mid&\mid\\ 1&\mathbf{X}&\mathbf{X}^{P_{2}}&\mathbf{X}^{P_{3}}&\cdots&\sin(\mathbf{X})&\cos(\mathbf{X})&\cdots\\ \mid&\mid&\mid&\mid&&\mid&\mid&\end{array}\right]$
where, higher polynomials are shown as $\mathbf{X}^{P_{2}}$, $\mathbf{X}^{P_{3}}$, $\cdots$. For example, $\mathbf{X}^{P_{2}}$ denotes the quadratic nonlinearities in the state $\mathbf{x}$. A potential function for the right-hand side of Eq 1 is represented by each column of $\mathbf{\Theta}(\mathbf{X})$ in Eq 3. The values of each column are often normalized to prevent numerical instability and ensure that features contribute equally to the model fitting process. Since only a small number of these terms are active in each row of $\mathbf{F}$, a sparse regression problem will be established to find the sparse vectors of coefficients $\mathbf{\Xi}=\left[\begin{array}[]{llll}\bm{\xi}_{1}&\bm{\xi}_{2}&\cdots&\bm{\xi}_{K}\end{array}\right]$ that determine which nonlinear terms are active:
$\dot{\mathbf{X}}=\mathbf{\Theta}(\mathbf{X})\mathbf{\Xi}$ (4)
@@
$\dot{\mathbf{x}}_{k}=\mathbf{F}_{k}\left(\mathbf{x}\right)=\mathbf{\Theta}({\mathbf{x}}^{T})\bm{\xi}_{k}$ (5)
@@
$\dot{\mathbf{x}}=\mathbf{F}\left(\mathbf{x}\right)=\mathbf{\Xi}^{T}\left(\mathbf{\Theta}({\mathbf{x}}^{T})\right)^{T}$ (6)
@@
$\min_{\mathbf{\Xi}}\frac{1}{2}\|\dot{\mathbf{X}}-\mathbf{\Theta}(\mathbf{X})\mathbf{\Xi}\|^{2}+\lambda\mathbf{R}(\mathbf{\Xi})$ (7)
where $\lambda$ is a hyper-parameter that controls the regularization’s strength and $\mathbf{R}(.)$ is a regularizer that imposes sparsity or other constraints on the solution. There are several well-known techniques for Eq 7 when $\mathbf{R}(.)$ is convex. The typical technique is to use $\mathbf{R}(.)$ to represent the sparsity-promoting $L1$ norm, which is a convex relaxation of the $L0$ norm. While LASSO [12] is commonly used to solve SINDy due to its convex formulation, it has several limitations such as bias in coefficient estimation. Hence, the SINDy algorithm employs sequential thresholded least squares (STLSQ), which given a parameter $\eta$ that specifies the minimum magnitude for a coefficient in $\mathbf{\Xi}$, performs a least squares fit and then zeros out all the coefficients with magnitude below the threshold. This process of fitting and thresholding is repeated until convergence. The SINDy algorithm has recently been integrated into a python package PySINDy [6].
@@
2.2 Graph Structured Time Series Data
Graphs provide a natural framework for representing systems where nodes (representing entities or states) and edges (representing interactions or connections) capture the underlying structure and dynamics of the network. The dynamics within a graph-structured system are governed by how changes in one node influence other nodes through their connections. This influence can be due to various factors such as physical interactions, information flow, or dependency relations. For instance, in a network of oscillators, the state of each oscillator at any given time is influenced by the states of the oscillators it is connected to, forming a complex web of interactions that evolve over time.
3 Materials and Method
3.1 Method
The SINDyG method is based on the SINDy algorithm (see Section 2). In SINDyG, shown in figure 1, the time histories of the state variables $\mathbf{X}$ are collected, their derivatives are calculated, and the structure of the network is identified. Next, a library of functions of the states, $\mathbf{\Theta}(\mathbf{X})$, is constructed. This feature library is used to find the fewest terms needed to satisfy $\dot{\mathbf{X}}=\mathbf{\Theta}(\mathbf{X})\mathbf{\Xi}$ while taking into account the network structure. The few active entries in the vectors of $\mathbf{\Xi}$ are identified using sparse regression [3]. Here, we provide details of each step of the SINDyG method.
3.1.1 Data Preparation
@@
3.1.2 Network Dynamics Discovery
@@
$\min_{\mathbf{\Xi}}\frac{1}{2}\|\dot{\mathbf{X}}-\mathbf{\Theta}(\mathbf{X})\mathbf{\Xi}\|_{2}^{2}+\lambda\|\mathbf{P}\odot\mathbf{\Xi}\|_{2}^{2}$ (8)
The parameter $\lambda$ controls the amount of regularization. The term $\mathbf{P}\odot\mathbf{\Xi}$ represents the element-wise product of $\mathbf{P}$ and $\mathbf{\Xi}$, while the penalty matrix $\mathbf{P}$ leverages the network’s adjacency matrix $\mathbf{A}$ to introduce a graph-aware penalty. The intuition behind the penalty matrix $\mathbf{P}$ is to determine the relative importance of different terms in the library based on the network’s connectivity. By encouraging coefficients linked to strongly connected nodes and discouraging those linked to weakly connected nodes, this penalty term guides the algorithm toward solutions that ensure the discovered dynamics are consistent with the underlying network topology.
In order to calculate this penalty, we iterate through each candidate term, $c$, in the library of functions. For each term, we identify the source indices, $S_{c}$, representing the state variables present in term $c$. We then determine the sink indices, $D_{c}$, which are the state variables reachable from $S_{c}$. For each sink index $k\in D_{c}$, we compute the mean connectivity, $m_{ck}$, between the source indices and sink index $k$ using the adjacency matrix $\mathbf{A}$. The penalty matrix is then calculated using a formula similar to the sigmoid formula:
$\mathbf{P}[c,k]=\frac{1}{1+\exp\left(\left(L/\left|S_{c}\right|\right)\cdot\left(m_{ck}-0.5\right)\right)}$ (9)
@@
3.1.3 Validating the Predicted Dynamics
Once the underlying equations governing a system have been identified using techniques such as SINDy or SINDyG, a comprehensive validation process is essential to assess the model’s performance. This involves evaluating the predicted model from multiple perspectives. Key metrics employed in this study include the model complexity index, Coefficient Error Index (CEI), generated signal accuracy, and training time.
@@ 3.1.3 Validating the Predicted Dynamics
The Coefficient Error Index (CEI) evaluates the accuracy of the estimated coefficients, reflecting the model’s fidelity to the underlying system. The CEI is calculated as the Mean Absolute Error (MAE) between the predicted coefficients and the true coefficients of the known dynamical model. Mathematically, the CEI is defined as:
$\text{CEI}=\frac{1}{KC}\sum_{k=1}^{K}\sum_{c=1}^{C}\left|\bm{\xi}_{k,c}-\widehat{\bm{\xi}}_{k,c}\right|,$ (10)
where $K$ is the number of dynamical equations, $C$ represents the number of possible candidate functions for each equation, $\bm{\xi}_{k,c}$ denotes the $c^{th}$ predicted coefficient for the $k^{th}$ equation, and $\widehat{\bm{\xi}}_{k,c}$ signifies the true coefficient. A lower CEI value indicates a closer match between the predicted and true coefficients, implying higher accuracy for the predicted dynamical model. Conversely, a higher CEI value suggests a larger discrepancy between the predicted and true values, signifying a potentially less accurate coefficient prediction. This method requires the availability of true coefficients. When the underlying governing equations are not readily available, alternative references such as values obtained from established physical laws or high-fidelity simulations can be employed.

### v3

@@ Abstract
The combination of machine learning (ML) and sparsity-promoting techniques is enabling direct extraction of governing equations from data, revolutionizing computational modeling in diverse fields of science and engineering. The discovered dynamical models could be used to address challenges in climate science, neuroscience, ecology, finance, epidemiology, and beyond. However, most existing sparse identification methods for discovering dynamical systems treat the whole system as one without considering the interactions between subsystems. As a result, such models are not able to capture small changes in the emergent system behavior. To address this issue, we developed a new method called Sparse Identification of Nonlinear Dynamical Systems from Graph-structured data (SINDyG), which incorporates the network structure into sparse regression to identify model parameters that explain the underlying network dynamics. We showcase the application of our proposed method using several case studies of neuronal dynamics, where we model the macroscopic oscillation of a population of neurons using the extended Stuart-Landau (SL) equation and utilize the SINDyG method to identify the underlying nonlinear dynamics. Our extensive computational experiments validate the improved accuracy and simplicity of discovered network dynamics when compared to the original SINDy approach. Sparse Regression, Model Extraction, Network Dynamics, Neural Dynamics, Stuart-Landau Model
@@ 1 Introduction
While SINDy as a general framework could be applied to a wide range of systems including networks  [22, 26, 8], it typically does not incorporate prior knowledge of the network structure during model identification. Instead, it treats all candidate terms equally, regardless of whether they reflect actual node-to-node connections or not. As a result, when applied to systems with many interacting components such as biological neural networks in the brain [17], power grids [20], social networks [1], and traffic systems [28], SINDy may lead to the selection of spurious terms that do not represent known connectivity patterns. This could potentially yield to biased and less arcuate models that do not reflect the true dynamics of the system.
In this paper, we propose an extension of the SINDy method, termed SINDyG (Sparse Identification of Nonlinear Dynamics for Graph-structured data), which explicitly accounts for the network structure during sparse regression. By incorporating the connections between nodes, our approach uncovers not only the individual dynamics of each node but also the mechanisms by which they interact, leading to more accurate and interpretable models. SINDyG introduces a graph-informed regularization strategy that uses prior knowledge of the network’s adjacency structure to guide model discovery. This enables more accurate and parsimonious identification of governing equations, particularly in systems where node-to-node interactions are sparse and structured. We test our method using several case studies of neuronal dynamics, where the goal is to identify nonlinear interactions between populations of neurons. Our results indicate improvements to model accuracy and simplicity when compared to the original SINDy method.
The rest of the work is organized as follows: In Sec. II, we discuss the necessary background information such as Sparse Identification of Nonlinear Dynamics (SINDy) [3] and graph-structured data. In Sec. III, we explain our proposed method for data-driven modeling of complex and interconnected dynamical systems. Next, Section IV presents the results of our method using several case studies based on neuronal dynamics. Finally, we conclude the paper in Section V.
@@ 2 Preliminaries
The following sections provide a description of Graph-structured time series data and Sparse Identification of Nonlinear Dynamics (SINDy), which are going to be frequently used in this paper.
2.1 Graph Structured Time Series Data
Graphs provide a natural framework for representing systems where nodes (representing entities or states) and edges (representing interactions or connections) capture the underlying structure and dynamics of the network. Graph-structured data refers to multivariate time series where each variable corresponds to a node in a graph, and the evolution of a node depends not only on its own state but also on the states of its connected neighbors. The dynamics within a graph-structured system are governed by how changes in one node influence other nodes through their connections. This influence can be due to various factors such as physical interactions, information flow, or dependency relations. For instance, in a network of oscillators, the state of each oscillator at any given time is influenced by the states of the oscillators it is connected to, forming a complex web of interactions that evolve over time.
In contrast to generic dynamical systems, where all candidate interactions are considered equally likely, graph-structured data imposes a prior structure on variable dependencies. In standard SINDy, the sparse regression step does not account for known network topology, often leading to the inclusion of spurious interactions. In this work, we incorporate graph structure directly into the regression process via a connectivity-aware penalty matrix.
@@ 2.2 Sparse Identification of Nonlinear Dynamics (SINDy)
$\frac{\mathrm{d}\mathbf{x}(t)}{\mathrm{d}t}=\mathbf{F}\left(\mathbf{x}(t)\right).$ (1)
@@
In order to compute the function $\mathbf{F}$ from data, a time series of system’s state vector $\mathbf{x}(t)$, which contains all the state variables, is collected. The derivatives $\mathbf{\dot{x}}(t)$ are then measured or numerically approximated and arranged in the following matrix:
[…]
@@
[…]
where, higher polynomials are shown as $\mathbf{X}^{P_{2}}$, $\mathbf{X}^{P_{3}}$, $\cdots$. For example, $\mathbf{X}^{P_{2}}$ denotes the quadratic nonlinearities in the state variables $\mathbf{x}$. A potential function for the right-hand side of Eq 1 is represented by each column of $\mathbf{\Theta}(\mathbf{X})$ in Eq 3. The values of each column are often normalized to prevent numerical instability and ensure that features contribute equally to the model fitting process. Since only a small number of these terms are active in each row of $\mathbf{F}$, a sparse regression problem will be established to find the sparse vectors of coefficients $\mathbf{\Xi}=\left[\begin{array}[]{llll}\bm{\xi}_{1}&\bm{\xi}_{2}&\cdots&\bm{\xi}_{K}\end{array}\right]$ that determine which nonlinear terms are active:
$\dot{\mathbf{X}}=\mathbf{\Theta}(\mathbf{X})\mathbf{\Xi}.$ (4)
@@
$\dot{\mathbf{x}}_{k}=\mathbf{F}_{k}\left(\mathbf{x}\right)=\mathbf{\Theta}({\mathbf{x}}^{T})\bm{\xi}_{k}.$ (5)
@@
$\dot{\mathbf{x}}=\mathbf{F}\left(\mathbf{x}\right)=\mathbf{\Xi}^{T}\left(\mathbf{\Theta}({\mathbf{x}}^{T})\right)^{T}.$ (6)
@@
$\min_{\mathbf{\Xi}}\frac{1}{2}\|\dot{\mathbf{X}}-\mathbf{\Theta}(\mathbf{X})\mathbf{\Xi}\|^{2}+\lambda\mathbf{R}(\mathbf{\Xi}),$ (7)
where $\lambda$ is a hyper-parameter that controls the regularization’s strength and $\mathbf{R}(.)$ is a regularizer that imposes sparsity or other constraints on the solution. There are several well-known techniques for Eq 7 when $\mathbf{R}(.)$ is convex. The typical technique is to use $\mathbf{R}(.)$ to represent the sparsity-promoting $L1$ norm, which is a convex relaxation of the $L0$ norm. While LASSO [13] is commonly used to solve SINDy due to its convex formulation, it has several limitations such as bias in coefficient estimation. Hence, the SINDy algorithm employs sequential thresholded least squares (STLSQ), which given a parameter $\eta$ that specifies the minimum magnitude for a coefficient in $\mathbf{\Xi}$, performs a least squares fit and then zeros out all the coefficients with magnitude below the threshold. This process of fitting and thresholding is repeated until convergence. The convergence properties of the STLSQ algorithm have been discussed in [27], which shows that, under assumptions such as exact sparsity of the true model and a well-conditioned feature matrix, the procedure converges to a stable sparse solution. The SINDy algorithm has recently been integrated into a python package PySINDy [7].
3 Materials and Method
The SINDyG method is based on the SINDy algorithm that was introduced in the previous section. In SINDyG, shown in figure 1, the time histories of the state variables $\mathbf{X}$ are collected, their derivatives are calculated, and the structure of the network is identified. Next, a library of functions of the state variables, $\mathbf{\Theta}(\mathbf{X})$, is constructed. This feature library is used to find the fewest terms needed to satisfy $\dot{\mathbf{X}}=\mathbf{\Theta}(\mathbf{X})\mathbf{\Xi}$ while taking into account the network structure. The few active entries in the vectors of $\mathbf{\Xi}$ are identified using sparse regression [3]. Here, we provide details of each step of the SINDyG method.
@@ 3 Materials and Method
3.1 Data Preparation
@@
3.2 Network Dynamics Discovery
@@
$\min_{\mathbf{\Xi}}\frac{1}{2}\|\dot{\mathbf{X}}-\mathbf{\Theta}(\mathbf{X})\mathbf{\Xi}\|_{2}^{2}+\lambda\|\mathbf{P}\odot\mathbf{\Xi}\|_{2}^{2}.$ (8)
The parameter $\lambda$ controls the amount of regularization. The term $\mathbf{P}\odot\mathbf{\Xi}$ represents the element-wise product of $\mathbf{P}$ and $\mathbf{\Xi}$, while the penalty matrix $\mathbf{P}$ leverages the network’s adjacency matrix $\mathbf{A}$ to introduce a graph-aware penalty. Unlike the original SINDy method, which applies uniform penalties to all terms in the candidate function library, SINDyG introduces a penalty matrix $\mathbf{P}$ that biases the regression toward selecting terms consistent with the known graph structure. The intuition behind the penalty matrix $\mathbf{P}$ is to determine the relative importance of different terms in the library based on the network’s connectivity. By encouraging coefficients linked to strongly connected nodes and discouraging those linked to weakly connected nodes, this penalty term guides the algorithm toward solutions that ensure the discovered dynamics are consistent with the underlying network topology.
In order to calculate this penalty, we iterate through each candidate term, $c$, in the library of functions. For each term, we identify the source indices, $S_{c}$, representing the state variables present in term $c$. We then determine the sink indices, $D_{c}$, which are the state variables reachable from $S_{c}$. For each sink index $k\in D_{c}$, we compute the mean connectivity, $m_{ck}$, between the source indices and sink index $k$ using the adjacency matrix $\mathbf{A}$. For a candidate term $c$ involving source variables $S_{c}$, we define the mean connectivity to sink node $k$ as:
$m_{ck}=\frac{1}{\left|S_{c}\right|}\sum_{i\in S_{c}}\mathbf{A}[i,k].$ (9)
This captures the average influence of the source variables on node $k$ under the network topology. The penalty matrix is then calculated using a formula similar to the sigmoid formula:
$\mathbf{P}[c,k]=\frac{1}{1+\exp\left(\left(L/\left|S_{c}\right|\right)\cdot\left(m_{ck}-0.5\right)\right)},$ (10)
@@
While Eq. 8 presents a ridge regression objective incorporating a graph-informed penalty matrix $\mathbf{P}$, sparsity is not enforced directly via $L0$ norm. Instead, we employ a modified version of the Sequentially Thresholded Least Squares (STLSQ) algorithm, which iteratively alternates between solving Eq. 8 and applying a hard threshold to small coefficients in $\mathbf{\Xi}$. This results in a sparse coefficient matrix over iterations. This iterative procedure ensures that the discovered model aligns with the known network structure while maintaining interpretability and computational efficiency. The complete graph-aware STLSQ procedure is summarized below.
Algorithm 1 Summary of Graph-aware STLSQ Procedure in SINDyG
0:  Time series data $X$, derivatives $\dot{X}$, adjacency matrix $A$, threshold $\eta$, penalty parameter $\lambda$
0:  Sparse coefficient matrix $\Xi$
1:  Construct the library matrix $\Theta(X)$
2:  Compute penalty matrix $P$ using mean connectivity based on $A$
3:  Initialize coefficient matrix $\Xi$
4:  repeat
5:     for each column $k$ of $\Xi$ do
6:        Scale $\Theta(X)$ and $\xi_{k}$ using penalty vector $p_{k}$ (column $k$ of $P$)
7:        Solve ridge regression subproblem:
$\min_{\xi_{k}^{\prime}}\|\dot{x}_{k}-\Theta^{\prime}(X)\xi_{k}^{\prime}\|_{2}^{2}+\lambda\|\xi_{k}^{\prime}\|_{2}^{2}$
8:        Revert scaling: $\xi_{k}\leftarrow\xi_{k}^{\prime}/p_{k}$
9:        Threshold small entries in $\xi_{k}$: set elements $<\eta$ to zero
10:     end for
11:  until convergence of $\Xi$
12:  return  $\Xi$
3.3 Validating the Predicted Dynamics
Once the underlying equations governing a system have been identified using techniques such as SINDy or SINDyG, a comprehensive validation process is essential to assess the model’s performance. This involves evaluating the predicted model from multiple perspectives. Key metrics employed in this study include the model complexity index, the coefficient Mean Absolute Error (MAE), the generated signal accuracy, and training time.
@@ 3.3 Validating the Predicted Dynamics
The coefficient mean absolute error (MAE) evaluates how closely the learned coefficients match the true coefficients. It is computed as the average absolute difference across all entries in $\bm{\xi}$. We chose MAE over mean squared error (MSE) to emphasize interpretability and robustness. MAE provides a clear indication of average deviation per term and is less influenced by outliers, which is desirable in sparse models where most terms are zero and a few incorrect activations can otherwise dominate the score. Mathematically, the coefficient MAE is defined as:
Coefficient $\text{Coefficient MAE}=\frac{1}{KC}\sum_{k=1}^{K}\sum_{c=1}^{C}\left|\bm{\xi}_{k,c}-\widehat{\bm{\xi}}_{k,c}\right|,$ (11)
where $K$ is the number of dynamical equations, $C$ represents the number of possible candidate functions for each equation, $\bm{\xi}_{k,c}$ denotes the $c^{th}$ predicted coefficient for the $k^{th}$ equation, and $\widehat{\bm{\xi}}_{k,c}$ signifies the true coefficient. A lower coefficient MAE value indicates a closer match between the predicted and true coefficients, implying higher accuracy for the predicted dynamical model. Conversely, a higher coefficient MAE value suggests a larger discrepancy between the predicted and true values, signifying a potentially less accurate coefficient prediction. This method requires the availability of true coefficients. When the underlying governing equations are not readily available, alternative references such as values obtained from established physical laws or high-fidelity simulations can be employed.
The model complexity index ($\gamma$) used throughout our experiments is directly tied to the sparsity of the coefficient matrix $\mathbf{\Xi}$, as it reflects the number of nonzero terms active in the discovered dynamical equations. A sparser matrix corresponds to a lower $\gamma$, indicating a simpler and more interpretable model. Similarly, the coefficient MAE is also indirectly linked to sparsity: a higher coefficient MAE often arises when additional spurious terms are incorrectly activated in $\mathbf{\Xi}$. This observation aligns with prior findings which noted that models with poor predictive performance tend to exhibit higher complexity and reduced sparsity [14]. By reporting both $\gamma$ and coefficient MAE, we capture how closely the learned dynamics match the ground truth in both structure and accuracy.